# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems

### v1

@@ Abstract.
Social media platforms are increasingly developing features that display crowdsourced context alongside posts, modeled after X’s Community Notes. These systems, which we term Crowdsourced Context Systems (CCS), have the potential to reshape our information ecosystem as major platforms embrace them as alternatives to top-down fact-checking. To deeply understand the features and implications of such systems, we perform a systematic literature review of existing CCS research and analyze several real-world CSS implementations. Based on our analysis, we develop a framework with three distinct components. First, we present a theoretical model to help conceptualize and define CCS. Second, we identify a design space encompassing six key aspects of CCS: participation, inputs, curation, presentation, platform treatment, and transparency. Third, we identify key normative implications of different CCS design and implementation choices. Our framework integrates these theoretical, design, and ethical perspectives to establish a foundation for future human-centered research on Crowdsourced Context Systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Social media platforms play an important role as sites of entertainment (Herder and Staring, 2024), news consumption (Chouaki et al., 2024), and interpersonal connection (Preece, 1999). At the same time, these platforms must contend with problems stemming from harmful user-uploaded content, which can take the form of abuse, spam, or misinformation (Scheuerman et al., 2021). Specifically, platforms have tried various forms of content moderation (Grimmelmann, 2015; Lampe and Resnick, 2004; Jhaver et al., 2019) to deal with misinformation, including centralized fact-checking efforts to label or remove it (Gillespie, 2018), as well as distributed approaches where users are empowered to tackle these challenges themselves (Seering, 2020). While these approaches have had some success, they have also evoked backlash and accusations of bias from both users and lawmakers, particularly on the political right (Jiang et al., 2020; Johnson, 2023). Recently, a new approach has emerged: beginning with X/Twitter, major social media platforms are increasingly introducing systems that crowdsource “helpful context” to display alongside potentially problematic posts (Coleman, 2021) (e.g. Figure 1). These systems, which we term Crowdsourced Context Systems (CCS), have seen growing adoption and impact, with some of the world’s largest platforms, including Meta (Facebook, Instagram, and Threads), TikTok, and YouTube all recently launching their own implementations. Meta has gone one step further, announcing that their CCS will replace their fact-checking program in the U.S. (Kaplan, 2025).
@@
As CCS adoption grows, these systems have the potential to reshape our information ecosystem. Early studies into such systems have produced promising results, finding that CCS can increase user informedness and decrease the spread of misinformation (Wojcik et al., 2022; Chuai et al., 2024b; Drolsbach and Pröllochs, 2023), although it is still unclear whether CCS can replace centralized fact-checking, as some platforms have done or may seek to do (Kaplan, 2025). As CCS develop, novel implementation questions also come to the fore, including how to design these systems most effectively, for example, to provide reliable context quickly enough to prevent the spread of the most viral content (Chuai et al., 2024b, a; Renault et al., 2024). At this critical moment of change in the information ecosystem, it is important to understand how CCS function, how we should evaluate them, and how these systems may be better designed.
@@
This paper makes two contributions towards the goal of better understanding CCS. Our first contribution is a systematic literature review of existing research studying implementations of CCS. Our review finds significant research into CCS contributor behavior and the impact of displaying notes. However, the review highlights the fact that certain aspects of these systems are underexplored by existing research. We thus set out to identify additional facets of these systems that may affect their impact. Our second contribution is an analysis and comparison of real-world implementations of CCS by several major social media platforms, in order to create a framework to guide future HCI research on these novel and important systems. Our framework is composed of three parts. First, we offer a theoretical model of a CCS that identifies these systems’ core characteristics and what sets them apart from prior crowdsourced systems. Our model posits that while CCS use familiar moderation tools like crowdsourced information and context labels, they are unique in how they are natively integrated into platforms with the explicit goal of fostering user informedness. Second, we outline a design space (MacLean et al., 1996) for CCS in order to taxonomize the important design decisions made in their implementations. Design spaces are a conceptual tool used in HCI research that can guide work on complex socioetechnical systems by structuring the space of design possibilities (Zhang et al., 2024; Lee et al., 2024; Card et al., 1990). We group these design decisions into six high-level aspects of CCS that combine and interact to shape their impact: participation, inputs, curation, platform treatment, and transparency. Third, we identify key normative implications of these systems in order to make sense of the ethical stakes involved in their design. As systems with the power to shape the information ecosystem, we emphasize three interconnected concerns: user informedness, distribution of power, and fairness. We discuss the role of CCS within the broader conversation about content moderation and establish a foundation for future human-centered research on Crowdsourced Context Systems.
@@ 2. Background
In January 2021, Twitter (now X) launched a U.S. pilot for the first CCS, which they called Birdwatch, presenting it as, “a community-based approach to misinformation” (Coleman, 2021). This system allowed a subset of users to suggest notes to, “provide informative context” about Tweets (posts) on the platform. The same subset of users then rated these suggested notes and these ratings were used as the inputs to a note curation algorithm that aimed to select notes that would be helpful to a broad audience. In this pilot, these notes were only visible on a separate site, but Twitter announced their plan, “to make notes visible directly on Tweets for the global Twitter audience, when there is consensus from a broad and diverse set of contributors.”
In March 2022, Twitter expanded the pilot, announcing that now, “a small (and randomized) group of people on Twitter in the US will see Birdwatch notes directly on some Tweets” (Coleman, 2022). Twitter was purchased by Elon Musk in October 2022, renamed X, and soon Birdwatch was renamed Community Notes. In December 2022, the pilot phase ended and users all over the world began seeing selected notes directly alongside Tweets (Community Notes [@CommunityNotes], 2022). Since then, the system has continued to evolve. A key feature of X’s Community Notes is its transparency: the note curation code is open source [^1] https://github.com/twitter/communitynotes/tree/main, as are the crowdsourced notes and ratings data collected by the system [^2] https://x.com/i/communitynotes/download-data.
In June 2024, YouTube announced a pilot for a similar system, which they called Notes (Team, 2024). In January 2025, Meta announced a pilot for their own implementation, also called Community Notes, which would launch on Facebook, Instagram, and Threads (Kaplan, 2025). Finally, TikTok announced their own pilot in April 2025, which they called Footnotes (Presser, [n. d.]). According to the platforms, all of these new systems are based on X’s open source implementation of the note curation algorithm. Unlike X’s implementation, though, none of the new systems are open source or open data. In fact, while all of these systems serve similar purposes, they differ from each other in crucial ways due to unique design decisions made by each platform. Since these systems are increasingly prominent features of some of the most important social media platforms, it is critical to define, understand, and reason about them so the HCI community can better inform their implementation and evaluate their impact.
@@ 3. Related Work
While CCS are a relatively new innovation, these systems relate to and touch on multiple bodies of work from the HCI literature, including: research on discourse and deliberation systems, research on fact-checking infrastructures, and research on crowdsourced responses to misinformation. We describe these areas here; later in the paper (Section 6) we provide a detailed review of the literature that is specifically addressing CCS.
@@
In a democracy that values free speech, public participation in decision-making via deliberation is considered a social good (Small, 2021). One classic concern of HCI research has thus been how to design systems that can support public deliberation among users with diverse, possibly conflicting, perspectives (Kriplean et al., 2012). One such approach was recently termed bridging systems, or systems that, “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (Ovadya and Thorburn, 2023). Such systems can be built with various goals in mind, such as finding common ground between peace-builders in armed conflict (Konya et al., 2025), or facilitating participatory approaches to allocate public funding (Yang and Bachmann, 2025). CCS can be viewed as highly distributed deliberation systems that aim to allow more informed public participation in social media. Community Notes, X’s CCS implementation, is one such system. Indeed, according to its creators, Community Notes was designed to identify crowdsourced content (notes), “that both inform understanding (decrease propensity to agree with a potentially misleading claim) and are seen as helpful by a diverse population of users” (Wojcik et al., 2022). CCS thus reflect a long-standing HCI concern with designing systems that support civic deliberation, adapting these ideas to meet the scale, speed, and additional challenges of major social media platforms.
@@
CCS also align with another classic goal of HCI research, which is to help users assesses the veracity of information that they encounter on social media. Emerging from journalistic practice, one approach to address this issue is to rely on external fact-checking (Walter et al., 2020; Young et al., 2018), where independent organizations verify public claims and publish their determinations. While these techniques can have positive effects, they are time-consuming and traditionally require a large amount of professional labor (Hassan et al., 2015; Juneja and Mitra, 2022), and thus may struggle keep up with the speed at which content is published on the internet (Micallef et al., 2022). As a result, researchers have explored more scalable approaches, such as automated tools to support human fact-checkers (Warren et al., 2025), as well as fully-automated approaches (Guo et al., 2022; Hassan et al., [n. d.]). While external fact-checks are often found on websites different from where the claim originated, such as Snopes [^3] https://www.snopes.com/ or Politifact [^4] https://www.politifact.com/, social media platforms may also provide their own fact-checks, which can be presented directly alongside user-posted content (How, [n. d.]). Studies have investigated both the most effective ways to present these fact-checks in order to shift belief (Clayton et al., 2020) and also the psychological characteristics that make some users reluctant to engage with them (Tanaka et al., 2023). In addition to issues of efficacy, and the large cost of running such fact-check programs, social media platforms have struggled with allegations and perceptions that their fact-checks are biased against certain political ideologies (Cotter et al., 2022).
While fact-check systems traditionally rely on a central authority, alternative approaches to dealing with information of questionable quality make use of the distributed expertise of a crowd (Pennycook and Rand, 2019; He et al., 2025). Crowd-based fact-checking is one such approach, which despite being cheaper and faster than professional fact-checking, has often been found to be similarly accurate and effective (Allen et al., 2021; Martel et al., 2024; Zhao and Naaman, 2023). HCI research has explored the practices of those who participate in crowdbased fact-checking (He and He, 2022), their potential biases (Draws et al., 2022), and how fact-checking tasks may be best distributed to a crowd (Bhuiyan et al., 2020). Beyond fact-checking, there are other crowd-based approaches to handling misinformation, such as digital juries (Fan and Zhang, 2020; Zhao and Hobbs, 2025), distributed moderation through up/down votes (Lampe and Resnick, 2004; Epstein et al., 2020), and community self-governance (Seering, 2020). CCS share much in common with these previous crowdsourced fact-checking systems, but they have a novel set of processes and structures that make CCS stand out as a new type of interactive systems. Informed by these past bodies of research, we next define what makes these new systems unique.
@@
4. Generalized Model of Crowdsourced Context Systems
We coin the term Crowdsourced Context System to describe these systems, a novel and important addition to the content moderation toolbox. Such systems have several core characteristics. First, CCS rely on contributions crowdsourced from users of the platform. Second, these systems are designed to specifically elicit context from the crowd in order to help users understand and reflect on the veracity of the original post (as opposed to other types of responses, like, say, the most humorous reply). Third, these systems supplement user content by displaying additional text, or counterspeech (Peng and Grimmelmann, 2024), alongside the original posts. Finally, these systems are natively integrated into social media platforms. Unlike organic crowd behavior that emerges among users, these systems are built and operated by the platforms, giving platform designers the power to influence how the systems function.
@@ 5. Methods
To better understand CCS and their design possibilities, we performed two distinct reviews. First, we performed a systematic literature review of existing research about CCS. Second, we performed inductive thematic analysis and affinity diagramming on a set of CCS design features extracted from documents describing the real-world CCS implementations of several major social media platforms.
@@ 5.1. Systematic Literature Review
Records identified through database searching (n = 205) Duplicates excluded (n = 8) Records after removing duplicates (n = 197) Records excluded due to title (n = 39) Records after title screen (n = 158) Records excluded due to abstract (n = 124) Records after abstract screen (n = 34) Reasons for full-text exclusion (n = 4) RQs are not about a CCS (n = 4) Studies included in qualitative synthesis (n = 30) IdentificationScreeningEligibilityIncluded
Figure 3. PRISMA flow diagram (Page et al., 2021) for our systematic literature review.:
In order to synthesize the research community’s understanding of Crowdsourced Context Systems, we performed a PRISMA-style systematic literature review (Page et al., 2021). Figure 3 shows a flow diagram summarizing our literature search and filtering process. Our goal was to identify original research that directly studies CCS. Since Twitter’s implementation was the first and most well-known CCS, we specifically used “Birdwatch” and “Community Notes” as keywords, as well as “Twitter” and “X”, with the expectation that any emerging systems would reference them. We additionally restricted our search to articles that have been written since 2021, when Birdwatch first launched. Note that since this is an emerging technology and the set of articles is still limited, we decided to include sources of knowledge that are valued in the HCI community but are sometimes excluded from literature reviews, such as pre-prints and non-full-length conference submissions. We thus settled on the following inclusion and exclusion criteria:
Inclusion criteria: the study is about Birdwatch or Community Notes or a similar system, either “in the wild” or in a controlled setting.
Exclusion criteria: The article was not written in English, does not present original research, is only an abstract/a summary of other research, or is not about the functioning of a CCS.
@@
We conducted our database search in July 2025. We searched seven databases that are well-known in the fields of HCI and social media research. We began with six that were used in another recent HCI literature review of social media research (Zhang et al., 2023): ACM Digital Library, IEEE Xplore, SAGE Journals, ScienceDirect (Elsevier), Taylor & Francis Online (Tandfonline), Emerald Insight. We additionally searched arXiv in order to also include preprints, which are common in these fields.
@@
("community notes" OR "birdwatch") AND ("x" OR "twitter")
This search resulted in n=205 results.
@@
We screened the search results by first considering their title. One researcher read the title for each result and removed articles that were either duplicates or were obviously off-topic, such as those that were not about social media systems. This resulted in n=158 articles.
@@
One researcher read the abstracts of the remaining articles to further screen the results. We removed redundant articles that only summarized research that was contained elsewhere in our sample, such as books and book chapters, editorials, abstracts, keynotes, and complete conference proceedings. We also removed articles that were topically irrelevant, such as those that did not focus on CCS, or that used data from a CCS, but only to answer research questions unrelated to the system. This resulted in n=34 articles.
@@
Two researchers then read the full-text of the remaining 34 articles and applied the inclusion and exclusion criteria in order to produce the final set of papers. The researchers met during the process to compare their results and discuss any differences or challenging cases until they were in agreement. This produced the final set of n=30 articles.
@@
Next the two researchers independently grouped like-papers into affinity groups (Beyer and Holtzblatt, 1999) based on the focus of the study. The researchers reconvened, compared their groups, and discussed the reasons for their groupings. Several papers were shifted between groupings based on the discussion. Once the researchers agreed upon the groupings, they came up with labels to describe each group.

### v2

@@ Abstract.
Social media platforms are increasingly developing features that display crowdsourced context alongside posts, modeled after X’s Community Notes. These systems, which we term Crowdsourced Context Systems (CCS), have the potential to reshape our information ecosystem as major platforms embrace them as alternatives to top-down fact-checking. To deeply understand the features and implications of such systems, we perform a systematic literature review of existing CCS research and analyze several real-world CCS implementations. Based on our analysis, we develop a framework with two distinct components. First, we present a theoretical model to help conceptualize and define CCS. Second, we identify a design space encompassing six key aspects of CCS: participation, inputs, curation, presentation, platform treatment, and transparency. We discuss key normative implications of different CCS design and implementation choices. Our paper integrates these theoretical, design, and ethical perspectives to establish a foundation for future human-centered research on Crowdsourced Context Systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Social media platforms play an important role as sites of entertainment (herderDigitalJunkfoodSocial2024), news consumption (chouakiWhatNewsPeople2024a), and interpersonal connection (preeceEmpathicCommunitiesBalancing1999). At the same time, these platforms must contend with problems stemming from harmful user-uploaded content, which can take the form of abuse, spam, or misinformation (scheuermanFrameworkSeverityHarmful2021). Specifically, platforms have tried various forms of content moderation (grimmelmann2015virtues; lampeSlashdotBurnDistributed2004; jhaverHumanMachineCollaborationContent2019) to deal with misinformation, including centralized fact-checking efforts to label or remove it (gillespieCustodiansInternetPlatforms2018), as well as distributed approaches where users are empowered to tackle these challenges themselves (seeringReconsideringSelfModerationRole2020). While these approaches have had some success, they have also evoked backlash and accusations of bias from both users and lawmakers, particularly on the political right (jiangReasoningPoliticalBias2020; johnsonFactsAllegationsPolitical2023). Recently, a new approach has emerged: beginning with X/Twitter, major social media platforms are increasingly introducing systems that crowdsource “helpful context” to display alongside potentially problematic posts (IntroducingBirdwatchCommunitybaseda) (e.g. Figure 1). These systems, which we term Crowdsourced Context Systems (CCS), have seen growing adoption and impact, with some of the world’s largest platforms, including Meta (Facebook, Instagram, and Threads), TikTok, and YouTube all recently launching their own implementations. Meta has gone one step further, announcing that their CCS will replace their fact-checking program in the U.S. (heatheraMoreSpeechFewer2025).
@@
As CCS adoption grows, these systems have the potential to reshape our information ecosystem. Early studies into such systems have produced promising results, finding that CCS can increase user informedness and decrease the spread of misinformation (wojcikBirdwatchCrowdWisdom2022b; rollout21; diffusion21), although it is still unclear whether CCS can effectively replace centralized fact-checking, as some platforms have done or may seek to do (heatheraMoreSpeechFewer2025). As CCS develop, novel implementation questions also come to the fore, including how to design these systems most effectively—for example, to provide reliable context quickly enough to prevent the spread of the most viral content (rollout21; chuaiCommunitybasedFactcheckingReduces2024a; renaultCollaborativelyAddingContext2024a). At this critical moment of change in the information ecosystem, it is important to understand how CCS function, how we should evaluate them, and how these systems may be better designed.
@@
This paper makes two contributions towards the goal of better understanding CCS. Our first contribution is a systematic literature review of existing research studying CCS implementations. Our review finds significant research into CCS contributor behavior and the impact of displaying notes. However, the review highlights the fact that certain aspects of these systems are underexplored by existing research. We thus set out to identify additional facets of these systems that may affect their impact. Our second contribution is an analysis and comparison of real-world implementations of CCS by several major social media platforms, in order to create a framework to guide future HCI research on these novel and important systems. Our framework is composed of two parts. First, we offer a theoretical model of a CCS that identifies these systems’ core characteristics and what sets them apart from prior crowdsourced systems. Our model posits that while CCS use familiar moderation tools like crowdsourced information and context labels, they are unique in how they are natively integrated into platforms with the explicit goal of facilitating content interpretation. Second, we outline a design space (macleanQuestionsOptionsCriteria1996) for CCS in order to taxonomize the important design decisions made in their implementations. Design spaces are a conceptual tool used in HCI research that can guide work on complex socioetechnical systems by structuring the space of design possibilities (zhangFormFromDesignSpace2024a; leeDesignSpaceIntelligent2024; cardDesignSpaceInput1990). We group these design decisions into six high-level aspects of CCS that combine and interact to shape their impact: participation, inputs, curation, platform treatment, and transparency. Based on our framework, we discuss key normative implications of these systems in order to make sense of the ethical stakes involved in their design. As systems with the power to shape the information ecosystem, we emphasize three interconnected concerns: user informedness, distribution of power, and fairness. We situate CCS within the broader conversation about content moderation and establish a foundation for future human-centered research on Crowdsourced Context Systems.
@@ 2. Background
In January 2021, Twitter (now X) launched a U.S. pilot for the first large-scale CCS, which they called Birdwatch, presenting it as “a community-based approach to misinformation” (IntroducingBirdwatchCommunitybaseda). This system allowed a subset of users to suggest notes to “provide informative context” about Tweets (posts) on the platform. The same subset of users then rated these suggested notes and these ratings were used as the inputs to a note curation algorithm that aimed to select notes that would be helpful to a broad audience. In this pilot, these notes were only visible on a separate site, but Twitter announced their plan “to make notes visible directly on Tweets for the global Twitter audience, when there is consensus from a broad and diverse set of contributors.”
In March 2022, Twitter expanded the pilot, announcing that now “a small (and randomized) group of people on Twitter in the US will see Birdwatch notes directly on some Tweets” (BuildingBetterBirdwatch). Twitter was purchased by Elon Musk in October 2022, renamed X, and soon Birdwatch was renamed Community Notes. In December 2022, the pilot phase ended and users all over the world began seeing selected notes directly alongside Tweets (BeginningTodayCommunity2022). Since then, the system has continued to evolve. A key feature of X’s Community Notes is its transparency: the note curation code is open source [^1] https://github.com/twitter/communitynotes/tree/main, as are the crowdsourced notes and ratings data collected by the system [^2] https://x.com/i/communitynotes/download-data.
In June 2024, YouTube announced a pilot for a similar system, which they called Notes (TestingNewWays). In January 2025, Meta announced a pilot for their own implementation, also called Community Notes, which would launch on Facebook, Instagram, and Threads (heatheraMoreSpeechFewer2025). Finally, TikTok announced their own pilot in April 2025, which they called Footnotes (TestingNewFeature). According to the platforms, all of these new systems are based on X’s open source implementation of the note curation algorithm. Unlike X’s implementation, though, none of the new systems are open source or open data. In fact, while all of these systems serve similar purposes, they differ from each other in crucial ways due to unique design decisions made by each platform. Since these systems are increasingly prominent features of some of the most important social media platforms, it is critical to define, understand, and reason about them so the HCI community can better inform their implementation and evaluate their impact.
@@ 3. Related Work
While CCS are a relatively new innovation, these systems relate to and touch on multiple bodies of work from the HCI literature, including: research on content moderation in public discourse systems, research on fact-checking infrastructures, and research on crowdsourced responses to misinformation. We describe these areas in general terms here; later in the paper (Section 6) we provide a detailed review of the literature that is examining specific implementations of CCS.
@@
In a democracy that values free speech, public participation in civil discussion is considered a social good (smallPolisEscalarDeliberacion2021). One classic concern of HCI research has thus been how to design systems that can support public discussion among users with diverse, possibly conflicting, perspectives (kripleanSupportingReflectivePublic2012; wohnHowHandleOnline2017). CCS are an example of such a system. Like past public discussion systems, CCS provide a form of “content moderation,” which can be used to encourage certain types of participation and discourage others (grimmelmann2015virtues). Content moderation can take different approaches: it can be “top-down,” where platforms make moderation decisions (gillespieCustodiansInternetPlatforms2018; gillespieNotRecommendReduction2022), or “bottom-up,” where users do (seeringReconsideringSelfModerationRole2020; lloyd25; lloydAIRulesCharacterizing2025; fieslerRedditRulesCharacterizing2018). Additionally, moderation can be human-powered, or use varying degrees of automation to keep up with large volumes of content (vaccaroContestabilityContentModeration2021; gorwa20). Community Notes, X’s CCS implementation, incorporates all of these approaches: it uses a centralized algorithm to aggregate moderation feedback that is sourced from users. Community Notes is an example of a bridging system, or an automated system designed to, “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (ovadyaBridgingSystemsOpen2023). Indeed, according to its creators, Community Notes was designed to identify crowdsourced content (notes) “that both inform understanding (decrease propensity to agree with a potentially misleading claim) and are seen as helpful by a diverse population of users” (wojcikBirdwatchCrowdWisdom2022b). CCS thus reflect a long-standing HCI concern with designing systems that support civil discussion through content moderation, adapting these ideas to meet the scale and speed of major social media platforms.
@@
One moderation approach frequently studied in HCI research is the use of top-down fact-checking to deal with misinformation (walterFactCheckingMetaAnalysisWhat2020; youngFactCheckingEffectivenessFunction2018). Emerging from journalistic practice, top-down fact-checking approaches empower certain individuals or organizations to verify public claims and publish determinations about their accuracy. While these techniques can have positive effects, they are time-consuming and traditionally require a large amount of professional labor (hassanDetectingCheckworthyFactual2015; junejaHumanTechnologicalInfrastructures2022), and thus may struggle to keep up with the speed at which content is published on the internet (micallefTrueFalseStudying2022). As a result, researchers have explored automated ways to scale the fact-checking process (warrenShowMeWork2025; guoSurveyAutomatedFactChecking2022; hassanQuestAutomateFactChecking). Beyond research into fact-check production, studies have also investigated the effects that fact-checks have on users who encounter them (eckerPsychologicalDriversMisinformation2022). This work has explored both the most effective ways to present fact-checks in order to shift user belief (claytonRealSolutionsFake2020), as well as the psychological characteristics that make some users reluctant to engage with them (tanakaWhoDoesNot2023). In addition to issues of efficacy, and the large cost of running such fact-check programs, social media platforms have struggled with allegations and perceptions that their fact-checks are biased against certain political ideologies (cotterFactCheckingCrisisCOVID192022; jiangReasoningPoliticalBias2020).
CCS can be seen as an evolution in content moderation that aims to address the limitations of top-down fact-checking by empowering platform users. In doing so, these new systems build on a long history of HCI research that has explored ways of harnessing the distributed wisdom of a crowd to produce accurate information in networked settings (pennycookFightingMisinformationSocial2019; heSurveyRoleCrowds2025). This research has shown that crowds can perform various tasks, such as sourcing on-the-ground information (daileyJournalistsCrowdsourcerersResponding2014; agapieCrowdsourcingFieldCase2015), or fact-checking political claims (kripleanIntegratingOndemandFactchecking2014). Studies have also found that despite being cheaper and faster than professional fact-checking, crowd-based fact-checking can be similarly accurate and effective in certain circumstances  (allenScalingFactcheckingUsing2021; martelCrowdsCanEffectively2024a; zhaoInsightsComparativeStudy2023). At the same time, crowd-based fact-checking is not a silver bullet, and other studies have shown that the composition of the crowd can affect performance (godelModeratingMobEvaluating2021), and that users may respond negatively to “social correction” that comes from the crowd (mosleh21). Crowds that are representative of the general population may be biased and cannot accurately perform every task (drawsEffectsCrowdWorker2022). Researchers have studied techniques to mitigate these biases (spiekerDiversePerspectivesCan2023) and have explored other ways that crowds can play a useful role in moderating misleading information, such as via digital juries (fanDigitalJuriesCivicsOriented2020; zhaoEffectsNonEffectsSocial2025), or via distributed moderation with up/down votes (lampeSlashdotBurnDistributed2004; epsteinWillCrowdGame2020). CCS share much in common with previously explored crowdsourced fact-checking systems, but they have a novel set of processes and structures that lead us to define them as a new class of system. Informed by these past bodies of research, we next define what makes these new systems unique.
@@
4. Generalized Model of Crowdsourced Context Systems for Social Media
We coin the term Crowdsourced Context System to describe these social media systems, a novel and important addition to the content moderation toolbox. Such systems have several core characteristics. First, CCS rely on contributions crowdsourced from users of the platform. Second, these systems are designed to specifically elicit context from the crowd that is meant to help users interpret the original post (as opposed to other types of responses—for example, the most humorous reply). Crowdsourced context can facilitate interpretation by providing additional information that helps readers assess and interpret the veracity, potential motivations, or broader social or political context of the original post. Third, these systems supplement user content by displaying additional text, or counterspeech (pengRescuingCounterspeechBridgingBased2024a), alongside the original posts. Finally, these systems are natively integrated into social media platforms. Unlike organic crowd behavior that emerges among users, these systems are built and operated by the platforms, giving platform designers the power to influence how the systems function.
@@ 5. Methods
To better understand CCS and their design possibilities, we performed two distinct reviews. First, we performed a systematic literature review of existing research on the first large-scale CCS implementation: Twitter’s Community Notes. Second, we performed inductive thematic analysis and affinity diagramming on a set of CCS design features extracted from documents describing the real-world CCS implementations of several major social media platforms.
@@ 5.1. Systematic Literature Review
Figure 3. PRISMA flow diagram (pagePRISMA2020Statement2021) for our systematic literature review, showing both the database search and the backwards reference search. We analyzed a final set of papers that combined the results of these two searches (n=56).:
In order to synthesize the research community’s understanding of Crowdsourced Context Systems, we performed a PRISMA-style systematic literature review (pagePRISMA2020Statement2021). We sourced relevant papers using both a database keyword search and a backwards reference search (jalaliSystematicLiteratureStudies2012). Figure 3 shows a flow diagram summarizing our search and screening process. Our goal was to identify original research that directly studies CCS implementations. Since this category of system has not been consistently named prior to our work, we focused our review on studies of the first large-scale implementation of a CCS by a major social media platform: Twitter’s Community Notes/Birdwatch. We thus used “Birdwatch” and “Community Notes” as keywords and restricted our search to articles that have been written since 2021, when Twitter first launched their implementation. Note that since this is an emerging technology and the set of articles is limited, we decided to include sources of knowledge that are valued in the HCI community but are sometimes excluded from literature reviews, such as pre-prints and non-full-length conference submissions. We thus settled on the following inclusion and exclusion criteria:
Inclusion criteria: the study is about a Twitter-style CCS, whether “in the wild” or in a controlled setting.
Exclusion criteria: The article was not written in English, does not present original research, is only an abstract/a summary of other research, or is not about the functioning of a Twitter-style CCS.
@@
We conducted our database search in November 2025. We searched seven databases that are well-known in the fields of HCI and social media research. We began with six that were used in another recent HCI literature review of social media research (zhangWhatWeMean2023): ACM Digital Library, IEEE Xplore, SAGE Journals, ScienceDirect (Elsevier), Taylor & Francis Online (Tandfonline), Emerald Insight. We additionally searched arXiv in order to also include preprints, which are common in these fields.
@@
"community notes" OR "birdwatch"
This search resulted in n=590 results.
@@
We screened the search results by first considering their title. One researcher read the title for each result and removed articles that were either duplicates or were obviously off-topic, such as those that were not about social media systems. This resulted in n=254 articles.
@@
One researcher read the abstracts of the remaining articles to further screen the results. We removed redundant articles that only summarized research that was contained elsewhere in our sample, such as books and book chapters, editorials, abstracts, keynotes, and complete conference proceedings. We also removed articles that were topically irrelevant, such as those that did not focus on a Twitter-style CCS, or that used data from such a system, but only to answer research questions unrelated to the system. This resulted in n=47 articles.
@@
Two researchers then read the full-text of the remaining 47 articles and applied the inclusion and exclusion criteria to screen more thoroughly. The researchers met during the process to compare their results and discuss any differences or challenging cases until they were in agreement. This produced a set of n=40 articles.
Backwards Reference Search
The first author then used this set of 40 articles as the starting set for a backwards reference search and extracted all of the n=2,207 references included in these articles. We de-duplicated these references by paper title, which resulted in a set of n=1,385 unique papers. We then performed the same screening steps that we used in the forward search: one researcher first screened the titles of the cited papers, which resulted in n=175 papers, then screened this set by abstract, which resulted in n=23, then finally two researchers performed a full-text screen. At the end of this screening process we had identified n=16 relevant papers that were not found in our forward search. We combined the sets from the two searches and produced a final set of n=56 papers.
@@
Next the two researchers independently grouped similar papers into affinity groups (beyer99) based on the primary focus of the study. The researchers reconvened, compared their groups, and discussed the reasons for their groupings. Several papers were shifted between groups based on the discussion. Once the researchers agreed upon the groups, they came up with labels to describe each group based on the focus of the studies (see Table 1).